# On the Price of Decentralization in Decentralized Detection

### v1

@@ A-B Proof of Theorem 2
To find the optimal $\lambda$ in (16), we set the derivative over $\lambda$ to zero,
$\sum_{j=1}^{n}c\,\mathsf{E}_{P_{0}}\left[\log\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right]-\frac{\mathsf{E}_{P_{1}}\left[c\left(\log\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right)\left(\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right)^{\lambda c}\right]}{\mathsf{E}_{P_{1}}\left[\left(\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right)^{\lambda c}\right]}=0.$
Plug in $\lambda=c^{-1}$, we have
$\;\sum_{j=1}^{n}c\,\mathsf{E}_{P_{0}}\left[\log\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right]-\frac{c\,\mathsf{E}_{P_{1}}\left[\left(\log\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right)\left(\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right)^{\lambda c}\right]}{\mathsf{E}_{P_{1}}\left[\left(\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right)^{\lambda c}\right]}$
$=$ $\;\sum_{j=1}^{n}\mathsf{E}_{P_{0}}\left[c\log\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right]-\frac{\mathsf{E}_{P_{1}}\left[c\left(\log\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right)\left(\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right)\right]}{\mathsf{E}_{P_{1}}\left[\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right]}$
$=$ $\;\sum_{j=1}^{n}\mathsf{E}_{P_{0}}\left[c\log\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right]-\mathsf{E}_{P_{0}}\left[c\log\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right]$
[…]
Since the term in the supremum of (16) is concave in $\lambda$, thus $\lambda=c^{-1}$ must be the optimal solution of $\lambda$ if we choose $r_{i}=c/\pi_{i}$ for all $i\in[n]$. Denote the choice of the $r$ as $\pi^{-1}$, we have
$\liminf_{t\rightarrow\infty}-\frac{1}{t}\log\beta_{i}^{(t)*}(\pi^{-1},\epsilon)\geq\sum_{j=1}^{n}\mathsf{E}_{P_{0}}\left[\log\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right]-\log\mathsf{E}_{P_{1}}\left[\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right]=\sum_{j=1}^{n}D_{\mathrm{KL}}\left(P_{j,0}\|P_{j,1}\right).$
Since the convergence rate in the decentralized regime must not outperform the rate in the centralized regime, we must have
@@ A-C Proof of Theorem 3
$\;\liminf_{t\rightarrow\infty}-\frac{1}{t}\log\alpha_{i}^{(t)}(r;\eta^{(t)}_{i},\gamma^{(t)}_{i})\geq\liminf_{t\rightarrow\infty}-\frac{1}{t}\log P_{0}\Bigg{\{}\sum_{j=1}^{n}Z_{j}^{(t)}\geq\gamma_{i}^{(t)}\Bigg{\}}\geq\inf_{x\in\mathcal{F}}\Lambda^{*}(x)$
$=$ $\;\inf_{x\in\mathcal{F}}\sup_{\lambda\in\mathbb{R}^{n}}\big{\{}\langle\lambda,x\rangle-\Lambda(\lambda)\big{\}}=\sup_{\lambda\geq 0}\Bigg{\{}\lambda\gamma_{i}^{(t)}-\sum_{j=1}^{n}\log\mathsf{E}_{P_{0}}\Bigg{[}\exp\Bigg{(}\lambda\pi_{j}r_{j}\log\frac{P_{j,1}(X_{j})}{P_{j,0}(X_{j})}\Bigg{)}\Bigg{]}\Bigg{\}}.$
@@
$\liminf_{t\rightarrow\infty}-\frac{1}{t}\log\alpha_{i}^{(t)}(\pi^{-1};\eta^{(t)}_{i},0)\geq$ $\;\sup_{\lambda\geq 0}\Bigg{\{}-\sum_{j=1}^{n}\log\mathsf{E}_{P_{0}}\Bigg{[}\exp\Bigg{(}\lambda\log\frac{P_{j,1}(X_{j})}{P_{j,0}(X_{j})}\Bigg{)}\Bigg{]}\Bigg{\}}$
@@ A-D Proof of Theorem 4
for all $i\in[n]$.
Furthermore, let $S_{n}^{2}$ and $\alpha_{n}$ denote the second and third central moment of
@@
where $X_{j}\sim P_{j,0}$ for all $j\in[n]$.
Let $\gamma_{i}^{(t)}$ be the threshold such that the type-I error at time $t$ is less than or equal to $\epsilon$, then
$\;\beta_{i}^{(t)}(r^{*};\eta^{(t)}_{i},\gamma^{(t)}_{i})\leq P_{1}\left\{\ell^{(t)}_{i}\leq\gamma_{i}^{(t)}\right\}$
@@
$=$ $\;\sum_{x\in\mathcal{X}^{t}}P_{0}^{\otimes t}(x)\exp\left\{\log\frac{P_{1}^{\otimes t}(x)}{P_{0}^{\otimes t}(x)}\right\}\mathbbm{1}\left\{\sum_{\tau=1}^{t}\sum_{j=1}^{n}\frac{\left[W^{\tau}\right]_{ij}}{\pi_{j}}\log\frac{P_{j,1}(x_{j}^{(\tau)})}{P_{j,0}(x_{j}^{(\tau)})}\leq\gamma_{i}^{(t)}\right\}$
@@
$=$ $\;\sum_{x\in\mathcal{X}^{t}}P_{0}^{\otimes t}(x)\exp\left\{\tilde{h}^{(t)}_{i}(x)+\varepsilon_{i}^{(t)}(x)\right\}\mathbbm{1}\left\{\tilde{h}_{i}^{(t)}(x)\leq\gamma_{i}^{(t)}\right\}$ (17)
@@
for all $x\in\mathcal{X}^{t}$.
Let
$\tilde{S}_{n}^{(t)}=\left(\frac{1}{t}\sum_{\tau=1}^{t}\sum_{j=1}^{n}\mathsf{Var}_{P_{0}}\left[\frac{\left[W^{\tau}\right]_{ij}}{\pi_{j}}\log\frac{P_{j,1}(x_{j})}{P_{j,0}(x_{j})}\right]\right)^{\frac{1}{2}}.$
@@
$\leq$ $\;\sum_{x\in\mathcal{X}^{t}}P_{0}^{\otimes t}(x)\exp\left\{\tilde{h}^{(t)}_{i}(x)+\varepsilon_{i}^{(t)}(x)\right\}\mathbbm{1}\left\{\underbrace{\frac{\tilde{h}_{i}^{(t)}(x)-\tilde{H}_{i}^{(t)}t}{\tilde{S}_{n}^{(t)}\sqrt{t}}}_{y^{(t)}}\leq\underbrace{\frac{\gamma_{i}^{(t)}-\tilde{H}_{i}^{(t)}t}{\tilde{S}_{n}^{(t)}\sqrt{t}}}_{\lambda^{(t)}}\right\}$
$\leq$ $\;\exp\left\{\frac{\rho}{1-\rho}\sqrt{\left(\frac{\pi_{i}}{1-\pi_{i}}\right)\sum_{j=1}^{n}\frac{1}{\pi_{j}}L_{j}^{2}}\right\}\sum_{y^{(t)}}P_{Y^{(t)}}(y^{(t)})\exp\left\{y^{(t)}\tilde{S}_{n}^{(t)}\sqrt{t}+\tilde{H}_{i}^{(t)}t\right\}\mathbbm{1}\left\{y^{(t)}\leq\lambda^{(t)}\right\}$
$=$ $\;\exp\left\{\tilde{H}_{i}^{(t)}t+\lambda^{(t)}\tilde{S}_{n}^{(t)}\sqrt{t}+\frac{\rho}{1-\rho}\sqrt{\left(\frac{\pi_{i}}{1-\pi_{i}}\right)\sum_{j=1}^{n}\frac{1}{\pi_{j}}L_{j}^{2}}\right\}\sum_{y^{(t)}\leq\lambda^{(t)}}P_{Y^{(t)}}(y^{(t)})\exp\left\{\underbrace{\left(y^{(t)}-\lambda^{(t)}\right)\tilde{S}_{n}^{(t)}\sqrt{t}}_{z}\right\}$
$=$ $\;\exp\left\{\tilde{H}_{i}^{(t)}t+\lambda^{(t)}\tilde{S}_{n}^{(t)}\sqrt{t}+\frac{\rho}{1-\rho}\sqrt{\left(\frac{\pi_{i}}{1-\pi_{i}}\right)\sum_{j=1}^{n}\frac{1}{\pi_{j}}L_{j}^{2}}\right\}\sum_{z\leq 0}P_{Y^{(t)}}\left(\frac{z}{\tilde{S}_{n}^{(t)}\sqrt{t}}+\lambda^{(t)}\right)e^{z}$
[…]
@@
Following the proof of Lemma 3 in [12] and [13], Lemma 3 could be extended to our case since each $\frac{[W^{\tau}]_{ij}}{\pi_{j}}\log\frac{P_{j,1}(X^{(\tau)}_{j})}{P_{j,1}(X^{(\tau)}_{j})}$ has its second and third moments in approximately the same scale with respect to $t$. In our case, the term $\sigma$ becomes the square of the mean of the $tn$ variance terms and $\alpha_{3}$ is the mean of the $tn$ third moment terms.
@@
$\Delta\lambda^{(t)}=\lambda^{(t)}-\lambda$
@@
$\Delta\Phi^{(t)}=$ $\;\Phi(\lambda^{(t)})-\Phi(\lambda^{(t)})-\frac{\tilde{\alpha}_{n}}{6\sqrt{2\pi t}(\tilde{S}_{n}^{(t)})^{3}}\left(1-(\lambda^{(t)})^{2}\right)e^{-\frac{1}{2}(\lambda^{(t)})^{2}}+o\left(\frac{1}{\sqrt{n}}\right),$
where $(\tilde{S}_{n}^{(t)})^{2}t,\tilde{\alpha}_{n}t$ are the second and third moment of
$\sum_{\tau=1}^{t}\sum_{j=1}^{n}\frac{\left[W^{\tau}\right]_{ij}}{\pi_{j}}\log\frac{P_{j,1}(X_{j})}{P_{j,0}(X_{j})}.$
If we let $\varphi_{n}(x)=\frac{\tilde{\alpha}_{n}}{6\tilde{S}_{n}^{2}}(1-x)^{2}$, we have
@@
$\Delta\Phi^{(t)}=$ $\;\Phi^{\prime}(\lambda)\Delta\lambda^{(t)}+o(\Delta\lambda^{(t)})=\Phi^{\prime}(\lambda)\Delta\lambda^{(t)}+o(\Delta\Phi^{(t)})$ (22)
@@
$(\tilde{S}_{n}^{(t)})^{2}-S_{n}^{2}=$ $\;\frac{1}{t}\sum_{\tau=1}^{t}\sum_{j=1}^{n}\left(\left(\frac{\left[W^{\tau}\right]_{ij}}{\pi_{j}}\right)^{2}-1\right)\mathrm{Var}_{P_{0}}\left[\log\frac{P_{j,1}(X_{j}^{(\tau)})}{P_{j,0}(X_{j}^{(\tau)})}\right]$
@@
$\tilde{S}_{n}^{(t)}-S_{n}\leq$ $\;\left(\frac{1}{S_{n}+\tilde{S}_{n}^{(t)}}\right)\frac{1}{t}\left(\frac{1+\pi_{\min}}{\pi_{\min}}\right)\left(\frac{\rho}{1-\rho}\right)\sqrt{\left(\frac{\pi_{i}}{1-\pi_{i}}\right)\sum_{j=1}^{n}\frac{1}{\pi_{j}}\sigma^{4}_{j}}$
$=$ $\;\left(\frac{1}{2S_{n}}+o(1)\right)\frac{1}{t}\left(\frac{1+\pi_{\min}}{\pi_{\min}}\right)\left(\frac{\rho}{1-\rho}\right)\sqrt{\left(\frac{\pi_{i}}{1-\pi_{i}}\right)\sum_{j=1}^{n}\frac{1}{\pi_{j}}\sigma^{4}_{j}}$
@@
$\beta_{i}^{(t)}(r^{*};\eta^{(t)}_{i},\gamma^{(t)}_{i})\leq$ $\;\exp\left\{H_{n}t+\lambda S_{n}\sqrt{t}-\varphi_{n}(\lambda)\right\}\exp\left\{\frac{\rho}{1-\rho}\sqrt{\frac{\pi_{i}}{1-\pi_{i}}}\sqrt{\sum_{j=1}^{n}\frac{1}{\pi_{j}}\left(\mathrm{D}\left(P_{j,0}\|P_{j,1}\right)\right)^{2}}\right\}$
×$\;\times\exp\left\{\frac{\rho}{1-\rho}\sqrt{\frac{\pi_{i}}{1-\pi_{i}}}\sqrt{\sum_{j=1}^{n}\frac{1}{\pi_{j}}L_{j}^{2}}\right\}\cdot A$ (26)
@@
We then deal with the remaining integral term.
Let
@@
$\beta^{(t)*}_{\mathrm{cen}}(\epsilon)=\exp\left\{H_{n}t+\lambda S_{n}\sqrt{t}-\frac{1}{2}\log t-\frac{1}{2}\log\left(2\pi\right)-\log S_{n}-\varphi_{n}(\lambda)-\frac{\lambda^{2}}{2}\right\}.$

### v2

@@ A-B Proof of Theorem 2
To find the optimal $\lambda$ in (19), we set the derivative over $\lambda$ to zero, that is,
00 0 $=\sum_{j=1}^{n}c\,\mathsf{E}_{P_{0}}\left[\log\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right]$
$\quad-\sum_{j=1}^{n}\frac{\mathsf{E}_{P_{1}}\left[c\left(\log\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right)\left(\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right)^{\lambda c}\right]}{\mathsf{E}_{P_{1}}\left[\left(\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right)^{\lambda c}\right]}.$
Plug in $\lambda=c^{-1}$, the above is satisfied because
$\frac{c\,\mathsf{E}_{P_{1}}\left[\left(\log\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right)\left(\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right)^{\lambda c}\right]}{\mathsf{E}_{P_{1}}\left[\left(\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right)^{\lambda c}\right]}$
$=\frac{\mathsf{E}_{P_{1}}\left[c\left(\log\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right)\left(\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right)\right]}{\mathsf{E}_{P_{1}}\left[\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right]}$
$=\mathsf{E}_{P_{0}}\left[c\log\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right].$
Since the term in the supremum of (19) is concave in $\lambda$, thus $\lambda=c^{-1}$ must be the optimal solution of $\lambda$ if we choose $r_{i}=c/\pi_{i}$ for all $i\in[n]$. Denote the choice of such an $r$ as $\pi^{-1}$, we have
$\liminf_{t\rightarrow\infty}-\frac{1}{t}\log\beta_{i}^{(t)*}(\pi^{-1},\epsilon)$
$\geq\sum_{j=1}^{n}\mathsf{E}_{P_{0}}\left[\log\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right]-\log\mathsf{E}_{P_{1}}\left[\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right]$
$=\sum_{j=1}^{n}D_{\mathrm{KL}}\left(P_{j,0}\|P_{j,1}\right).$
Since the convergence rate in the decentralized regime cannot outperform the rate in the centralized regime, we must have
@@ A-C Proof of Theorem 3
$\liminf_{t\rightarrow\infty}-\frac{1}{t}\log\alpha_{i}^{(t)}(r;\eta^{(t)}_{i},\gamma^{(t)}_{i})$
$\geq\liminf_{t\rightarrow\infty}-\frac{1}{t}\log P_{0}\Bigg{\{}\sum_{j=1}^{n}Z_{j}^{(t)}\geq\gamma_{i}^{(t)}\Bigg{\}}$
$\geq\inf_{x\in\mathcal{F}}\Lambda^{*}(x)$
$=\inf_{x\in\mathcal{F}}\sup_{\lambda\in\mathbb{R}^{n}}\big{\{}\langle\lambda,x\rangle-\Lambda(\lambda)\big{\}}$
$=\sup_{\lambda\geq 0}\Bigg{\{}\lambda\gamma_{i}^{(t)}-\sum_{j=1}^{n}\log\mathsf{E}_{P_{0}}\Bigg{[}\exp\Bigg{(}\lambda\pi_{j}r_{j}\log\frac{P_{j,1}(X_{j})}{P_{j,0}(X_{j})}\Bigg{)}\Bigg{]}\Bigg{\}}.$
@@
$\liminf_{t\rightarrow\infty}-\frac{1}{t}\log\alpha_{i}^{(t)}(\pi^{-1};\eta^{(t)}_{i},0)$
$\geq\sup_{\lambda\geq 0}\Bigg{\{}-\sum_{j=1}^{n}\log\mathsf{E}_{P_{0}}\Bigg{[}\exp\Bigg{(}\lambda\log\frac{P_{j,1}(X_{j})}{P_{j,0}(X_{j})}\Bigg{)}\Bigg{]}\Bigg{\}}$
@@ A-D Proof of Theorem 4
for all $i\in[n]$. Furthermore, let $S_{n}^{2}$ and $\alpha_{n}$ denote the second and third central moment of
@@
where $X_{j}\sim P_{j,0}$ for all $j\in[n]$. Let $\gamma_{i}^{(t)}$ be the threshold such that the type-I error at time $t$ is less than or equal to $\epsilon$, then
$\beta_{i}^{(t)}(r^{*};\eta^{(t)}_{i},\gamma^{(t)}_{i})$
$\leq P_{1}\left\{\ell^{(t)}_{i}\leq\gamma_{i}^{(t)}\right\}$
@@
$=\sum_{x\in\mathcal{X}^{t}}\left\{\begin{array}[]{l}P_{0}^{\otimes t}(x)\exp\left\{\log\frac{P_{1}^{\otimes t}(x)}{P_{0}^{\otimes t}(x)}\right\}\\ \mathbbm{1}\left\{\sum_{\tau=1}^{t}\sum_{j=1}^{n}\frac{\left[W^{\tau}\right]_{ij}}{\pi_{j}}\log\frac{P_{j,1}(x_{j}^{(\tau)})}{P_{j,0}(x_{j}^{(\tau)})}\leq\gamma_{i}^{(t)}\right\}\end{array}\right\}$ (22)
@@
$=\sum_{x\in\mathcal{X}^{t}}P_{0}^{\otimes t}(x)\exp\left\{\tilde{h}^{(t)}_{i}(x)+\varepsilon_{i}^{(t)}(x)\right\}\mathbbm{1}\left\{\tilde{h}_{i}^{(t)}(x)\leq\gamma_{i}^{(t)}\right\},$ (23)
@@
Proof:
The proof is similar to the one for Proposition 3 in [14]. We have
$\left(\sum_{j=1}^{n}\left|\left[W^{\tau}\right]_{ij}-\pi_{j}\right|r_{j}\right)^{2}$
$\leq\left(\sum_{j=1}^{n}\frac{1}{\pi_{j}}\left|\left[W^{\tau}\right]_{ij}-\pi_{j}\right|^{2}\right)\left(\sum_{j=1}^{n}\pi_{j}r_{j}^{2}\right)$ (Cauchy-Schwarz)
and
$\sum_{j=1}^{n}\frac{1}{\pi_{j}}\left|\left[W^{\tau}\right]_{ij}-\pi_{j}\right|^{2}$
$=\sum_{j=1}^{n}\left(\frac{1}{\pi_{j}}\left(\left[W^{\tau}\right]_{ij}\right)^{2}-2\left[W^{\tau}\right]_{ij}+\pi_{j}\right)$
Following the remaining part in the proof for Proposition 3 in [14], Lemma 2 is proved. ∎
@@
for all $x\in\mathcal{X}^{t}$. Let
$\tilde{S}_{n}^{(t)}$ $=\left(\frac{1}{t}\sum_{\tau=1}^{t}\sum_{j=1}^{n}\mathsf{Var}_{P_{0}}\left[\frac{\left[W^{\tau}\right]_{ij}}{\pi_{j}}\log\frac{P_{j,1}(x_{j})}{P_{j,0}(x_{j})}\right]\right)^{\frac{1}{2}},$
$y^{(t)}$ $=\frac{\tilde{h}_{i}^{(t)}(x)-\tilde{H}_{i}^{(t)}t}{\tilde{S}_{n}^{(t)}\sqrt{t}},\ \lambda^{(t)}=\frac{\gamma_{i}^{(t)}-\tilde{H}_{i}^{(t)}t}{\tilde{S}_{n}^{(t)}\sqrt{t}},$
@@
$\leq\sum_{x\in\mathcal{X}^{t}}P_{0}^{\otimes t}(x)\exp\left\{\tilde{h}^{(t)}_{i}(x)+\varepsilon_{i}^{(t)}(x)\right\}\mathbbm{1}\left\{y^{(t)}\leq\lambda^{(t)}\right\}$
$\leq\exp\left\{\frac{\rho}{1-\rho}\sqrt{\left(\frac{\pi_{i}}{1-\pi_{i}}\right)\sum_{j=1}^{n}\frac{1}{\pi_{j}}L_{j}^{2}}\right\}$
$\quad\cdot\sum_{y^{(t)}}P_{Y^{(t)}}(y^{(t)})e^{y^{(t)}\tilde{S}_{n}^{(t)}\sqrt{t}+\tilde{H}_{i}^{(t)}t}\mathbbm{1}\left\{y^{(t)}\leq\lambda^{(t)}\right\}$
$=e^{\tilde{H}_{i}^{(t)}t+\lambda^{(t)}\tilde{S}_{n}^{(t)}\sqrt{t}+\frac{\rho}{1-\rho}\sqrt{\left(\frac{\pi_{i}}{1-\pi_{i}}\right)\sum_{j=1}^{n}\frac{1}{\pi_{j}}L_{j}^{2}}}$
$\quad\cdot\sum_{y^{(t)}\leq\lambda^{(t)}}P_{Y^{(t)}}(y^{(t)})\exp\Big{\{}\underbrace{\left(y^{(t)}-\lambda^{(t)}\right)\tilde{S}_{n}^{(t)}\sqrt{t}}_{z}\Big{\}}$
$=e^{\tilde{H}_{i}^{(t)}t+\lambda^{(t)}\tilde{S}_{n}^{(t)}\sqrt{t}+\frac{\rho}{1-\rho}\sqrt{\left(\frac{\pi_{i}}{1-\pi_{i}}\right)\sum_{j=1}^{n}\frac{1}{\pi_{j}}L_{j}^{2}}}$
$\quad\cdot\sum_{z\leq 0}P_{Y^{(t)}}\left(\frac{z}{\tilde{S}_{n}^{(t)}\sqrt{t}}+\lambda^{(t)}\right)e^{z}$
$=e^{\tilde{H}_{i}^{(t)}t+\lambda^{(t)}\tilde{S}_{n}^{(t)}\sqrt{t}+\frac{\rho}{1-\rho}\sqrt{\left(\frac{\pi_{i}}{1-\pi_{i}}\right)\sum_{j=1}^{n}\frac{1}{\pi_{j}}L_{j}^{2}}}$
$\quad\cdot\int_{z\leq 0}e^{z}dF_{Y^{(t)}}\left(\frac{z}{\tilde{S}_{n}^{(t)}\sqrt{t}}+\lambda^{(t)}\right).$ (25)
@@
Following the proofs in [12] and [13], the above Lemma 3 could be extended to our case since each $\frac{[W^{\tau}]_{ij}}{\pi_{j}}\log\frac{P_{j,1}(X^{(\tau)}_{j})}{P_{j,1}(X^{(\tau)}_{j})}$ has its second and third moments in approximately the same scale with respect to $t$. In our case, the term $\sigma$ becomes the square of the mean of the $tn$ variance terms and $\alpha_{3}$ is the mean of the $tn$ third moment terms.
@@
$\Delta\lambda^{(t)}$ $=\lambda^{(t)}-\lambda,$
@@
$\Delta\Phi^{(t)}$ $=\Phi(\lambda^{(t)})-\Phi(\lambda^{(t)})$
$\quad-\frac{\tilde{\alpha}_{n}}{6\sqrt{2\pi t}(\tilde{S}_{n}^{(t)})^{3}}\left(1-(\lambda^{(t)})^{2}\right)e^{-\frac{1}{2}(\lambda^{(t)})^{2}}$
$\quad+o\left(1/\sqrt{n}\right),$
where $(\tilde{S}_{n}^{(t)})^{2}t$ and $\tilde{\alpha}_{n}t$ are the second and third moments of
$\sum_{\tau=1}^{t}\sum_{j=1}^{n}\frac{\left[W^{\tau}\right]_{ij}}{\pi_{j}}\log\frac{P_{j,1}(X_{j})}{P_{j,0}(X_{j})}$
respectively. If we let $\varphi_{n}(x)=\frac{\tilde{\alpha}_{n}}{6\tilde{S}_{n}^{2}}(1-x)^{2}$, we have
@@
$\Delta\Phi^{(t)}$ $=\Phi^{\prime}(\lambda)\Delta\lambda^{(t)}+o(\Delta\lambda^{(t)})=\Phi^{\prime}(\lambda)\Delta\lambda^{(t)}+o(\Delta\Phi^{(t)})$
@@
$(\tilde{S}_{n}^{(t)})^{2}-S_{n}^{2}$
$=\frac{1}{t}\sum_{\tau=1}^{t}\sum_{j=1}^{n}\left(\left(\frac{\left[W^{\tau}\right]_{ij}}{\pi_{j}}\right)^{2}-1\right)\mathrm{Var}_{P_{0}}\left[\log\frac{P_{j,1}(X_{j}^{(\tau)})}{P_{j,0}(X_{j}^{(\tau)})}\right]$
@@
$\tilde{S}_{n}^{(t)}-S_{n}$
$\leq\frac{\frac{1}{t}\left(\frac{1+\pi_{\min}}{\pi_{\min}}\right)\left(\frac{\rho}{1-\rho}\right)}{S_{n}+\tilde{S}_{n}^{(t)}}\sqrt{\left(\frac{\pi_{i}}{1-\pi_{i}}\right)\sum_{j=1}^{n}\frac{1}{\pi_{j}}\sigma^{4}_{j}}$
$=\left(\frac{1}{2S_{n}}+o(1)\right)\frac{\left(\frac{1+\pi_{\min}}{\pi_{\min}}\right)\left(\frac{\rho}{1-\rho}\right)}{t}\sqrt{\left(\frac{\pi_{i}}{1-\pi_{i}}\right)\sum_{j=1}^{n}\frac{1}{\pi_{j}}\sigma^{4}_{j}}$
@@
$\beta_{i}^{(t)}(r^{*};\eta^{(t)}_{i},\gamma^{(t)}_{i})$
$\leq\exp\left\{H_{n}t+\lambda S_{n}\sqrt{t}-\varphi_{n}(\lambda)\right\}$
$\quad\cdot\exp\left\{\frac{\rho}{1-\rho}\sqrt{\frac{\pi_{i}}{1-\pi_{i}}}\sqrt{\sum_{j=1}^{n}\frac{1}{\pi_{j}}\left(\mathrm{D}\left(P_{j,0}\|P_{j,1}\right)\right)^{2}}\right\}$
$\quad\cdot\exp\left\{\frac{\rho}{1-\rho}\sqrt{\frac{\pi_{i}}{1-\pi_{i}}}\sqrt{\sum_{j=1}^{n}\frac{1}{\pi_{j}}L_{j}^{2}}\right\}\cdot A$ (31)
@@
We then deal with the remaining integral term. Let
@@
[…]
$=\left(\frac{1}{S_{n}}+o(1)\right)\frac{1}{\sqrt{t}}\int_{z\leq 0}e^{z-\frac{\lambda}{2}+o(1)}+\frac{1}{\sqrt{t}}B(\lambda)$
[…]
@@
$\beta^{(t)*}_{\mathrm{cen}}(\epsilon)=e^{H_{n}t+\lambda S_{n}\sqrt{t}-\frac{1}{2}\log t-\frac{1}{2}\log\left(2\pi\right)-\log S_{n}-\varphi_{n}(\lambda)-\frac{\lambda^{2}}{2}}.$